Could someone please explain to me how I can solve these three questions?
t - 7 < -15
-8 - b ≤ - 8
c + 5.1 ≤ -2.4
Thank you very much for your help.

Found 3 solutions by Theo, Alan3354, dkppathak:
Answer by Theo(13342)

t - 7 < -15
add 7 to both sides of the inequality to get t < -15 + 7
combine like terms to get t < -8

-8 - b <= -8
add 8 to both sides of the inequality to get -b <= 0
multiply both sides of the inequality by -1 to get b >= 0

when you multiply or divide both sides of an inequality by a negative number, the inequality reverses.

c + 5.1 <= -2.4
subtract 5.1 from both sides of the inequality to get c <= -2.4 - 5.1
combine like terms to get c <= -7.5

you solve inequalities just like you solve equations with the exception that, if you multiply or divide the inequality by a negative number, the inequality reverses.

also, if your unknown variable is on the right and you want to make it on the left, then you need to reverse the inequality.

in your second problem, you could have solved it two ways.

the problem was -8 - b <= -8

i solved it above by adding 8 to both sides of the inequality to get -b <= -8 + 8 which resulted in -b <= 0.
i then multiplied both sides of the inequality by -1 to get b >= 0.

the other way to solve it would have been add b to both sides of the inequality to get -8 <= b - 8.
i then added 8 to both sides of the inequality to get 0 <= b.
to switch sides, i then needed to reverse the inequality to get b >= 0.

here's a reference on solving inequalities you might find useful.

https://www.mathsisfun.com/algebra/inequality-solving.html


Answer by Alan3354(69443)

Could someone please explain to me how I can solve these three questions?
t - 7 < -15
Add 7 to each side.
=====================
-8 - b ≤ - 8
Add b, then add 8
===================
c + 5.1 ≤ -2.4
Subtract 5.1
===========================
PS Those are not questions, they're inequalities.
and, questions are not solved. Questions are answered.

Answer by dkppathak(439)

t - 7 < -15
-8 - b ≤ - 8
c + 5.1 ≤ -2.4
first
t-7<-15 adding 7 both side
t<-8 answer
second
-8-b<-8 adding 8 both side
-b<0 than b>0 answer
third
c+5.1<-2.4 by subtracting 5.1 both side
c<-7.5 answer